# Graph Agent Network: Empowering Nodes with Decentralized Communications Capabilities for Adversarial Resilience

### v1

@@ V-C1 Symmetry of Attentions
We further illustrate the fluctuation range of bidirectional attention on the same edge to provide additional evidence for effective communication between agents and to validate the efficacy of the neighbor feature aggregator. We select the node with the largest degree (denoted as $v_{c}$) in the Cora dataset as the central node and observe the forward and reverse attention of the edges formed with its 72 randomly selected neighbors. As shown in Fig. 6, it can be observed that the overall trends of forward and backward attention during the training process are similar, accompanied by tolerable fluctuations, and both converge to approximately the same values.
V-C2 Distribution of Attentions
. Here we investigate the attention distribution learned by agents for both normal edges and adversarial edges. A significant distinction between the two would indicate that agents, through communication, have become aware that lower weights should be assigned to adversarial edges, thereby autonomously filtering out information from illegitimate neighbors introduced by such edges. As the previous set of experiments have demonstrated the symmetry of attention, we present the attention of a randomly selected agent on one end of the edge. The experimental results, as depicted in Fig. 7, reveal a notable difference in the kernel density estimation (KDE) of attentions between normal edges and adversarial edges (generated by Metattack). This outcome substantiates that agents, by training the aggregator, have acquired the capability to autonomously filter out adversarial edges.
V-D Effectiveness of Embedding
Upon completing communication, the agent utilizes a well-trained embedding function to embed its feature into the label space. We conduct a visualization of the global embeddings to investigate whether the agents have acquired effective embeddings. If the agents’ embeddings have not undergone global training and are solely trained based on their own labels under knowledge-limited conditions, yet still exhibit clustering phenomena globally, it demonstrates the effectiveness of the embedding function. Figure 8 displays the embedding of the Cora and Citeseer and visualized by t-SNE [55], where points of different colors represent distinct categories. As can be observed, points from different categories form clear clusters, validating the efficacy of the embedding function.
Figure 8: Well-trained agents generated embedding.
@@ V-D Effectiveness of Embedding
Additionally, We provide a detailed account of the training process of node embeddings by GAgN on the Cora dataset. The experimental results are presented in Fig. 9. As can be seen, initially, the embeddings of all nodes are randomly dispersed in the label space. Through communication among agents, the nodes gradually reach a consensus on the embeddings, and by continuously adjusting, they gradually classify themselves correctly and achieve a steady state.
V-E Effectiveness of Degree Inference
Each agent possesses the ability to infer the degree of any node based on the degree inference function. To validate the effectiveness of this inference, we select 150 representative nodes as test sets after training the GAgN model on the corresponding clean graph, following these rules: 1) the top 50 nodes with the highest degree, denoted as $\{L$-$nodes\}$; 2) the top 50 nodes with the lowest degree, denoted as $\{S$-$nodes\}$; 3) the 50 nodes furthest away from the inference node in terms of graph distance, denoted as $\{D$-$nodes\}$. Notably, to avoid training set leakage into the test set, an agent’s 1-hop neighbors and select 2-hop neighbors are excluded from its test set. The experimental results, as shown in Table II, demonstrate that even without data on the test nodes, agents can generalize their inference capabilities to the nodes of the graph under limited knowledge training, effectively inferring the degree of nodes.
TABLE II: Degree inference accuracy (%)
Test set $\{L$-$nodes\}$$\{S$-$nodes\}$$\{D$-$nodes\}$
Cora 90.35 95.31 92.64
Citeseer 89.21 94.19 91.08
Polblogs 92.41 96.32 92.57
V-F Effectiveness of Neighboring Confidence
To fully validate the effectiveness of the neighboring confidence function on graph $\mathcal{G}$, $|\mathcal{V}\times\mathcal{V}|$ matrix operations are required. In order to reduce the computational complexity while ensuring a fair evaluation process, we construct four representative graphs and utilize the neighbor relationships within these graphs as test samples to create a test set that encompasses potential distinct task attributes: 1) The original graph $\mathcal{G}_{0}$, with test sample labels all set to 1, 2) A graph $\mathcal{G}_{1}$ generated by randomly distributing edges over V, with test sample labels determined by whether the generated edges are original or not, 3) A randomly rewired graph $\mathcal{G}_{2}$, with test sample labels all set to 0, and 4) A graph $\mathcal{G}_{3}$ containing only adversarial edges, with test sample labels all set to 0. Upon constructing the test sets, we randomly select 50 nodes to compute the confidence for neighbor relationships in different test sets, treating this task as a binary classification problem. The classification accuracy represents the effectiveness of the neighboring confidence function. Figure 10 displays the classification accuracy on different test sets constructed within various graphs, demonstrating that the neighboring confidence function achieves satisfactory accuracy in all test sets, thus showcasing its effectiveness.
Figure 10: Accuracy of neighboring Confidence function.
V-G Ablation Experiments
We employed a multi-module filtering approach to reduce computational complexity when filtering adversarial edges. To investigate the individual effects of these modules when operating independently, it is crucial to conduct ablation experiments. Section V-C has explored the distribution patterns of attention scores for normal and adversarial edges. Edges with low attention scores are initially considered suspicious, which implies that the attention threshold may influence the detection rate and false alarm rate. To quantitatively examine the impact of threshold selection on these rates, we utilize receiver operating characteristic (ROC) curves. As depicted in Fig. 11, the detection rate is relatively insensitive to the attention threshold. This allows for a high detection rate while maintaining low false alarm rates.
Figure 11: ROC curves with different attention threshold.
TABLE III: Results of ablation experiments
Dataset Cora Citeseer
Metrics TPR FPR TPR FPR
Att. 85.52% 5.56% 82.66% 7.20%
Att.+$\mathcal{D}$+$\mathcal{N}$ 83.04% 0.87% 79.60% 1.16%
After identifying suspicious edges based on attention, the degree inference function and neighboring confidence function collaboratively eliminate normal edges from the suspicious ones, further reducing false positive rates. Here we investigate whether these two functions effectively pinpoint adversarial edges while preserving normal edges as much as possible. We utilize true positive rate (TPR) to quantify the detection rate of adversarial edges, and false positive rate (FPR) to measure the misclassification rate of normal edges. We conduct Metattack as the attack method. The experimental results are shown in Table III, where “Att.” denotes detection using only attention, and “Att.+$\mathcal{D}$+$\mathcal{N}$” represents the combined detection of all three methods. It can be observed that solely relying on attention for detection results in a higher detection rate but also introduces a higher false positive rate. By incorporating both the degree inference function and neighboring confidence function, the false positive rate is significantly reduced without substantially affecting the detection rate.
VI Conclusions

### v2

@@ B.1 Symmetry of Attention in the Communication Process
In Section 3, we demonstrate the symmetry of well-trained attention on the same edge. Here, we further illustrate the fluctuation range of bidirectional attention on the same edge to provide additional evidence for effective communication between agents and to validate the efficacy of the neighbor feature aggregator. We select the node with the largest degree (denoted as $v_{c}$) in the Cora dataset as the central node and observe the forward and reverse attention of the edges formed with its 72 randomly selected neighbors. As shown in Figure 10, it can be observed that the overall trends of forward and backward attention during the training process are similar, accompanied by tolerable fluctuations, and both converge to approximately the same values.
B.2 Embedding in the Communication Process
@@
We provide a detailed account of the training process of node embeddings by GAgN on the Cora dataset. The experimental results are presented in Figure 11. As can be seen, initially, the embeddings of all nodes are randomly dispersed in the label space. Through communication among agents, the nodes gradually reach a consensus on the embeddings, and by continuously adjusting, they gradually classify themselves correctly and achieve a steady state.